In Newton's ring experiment the diameter of 16 ^ (th) bright ring changes from 1.8 cm to 1.5cm when a liquid is introduced the plate and the lens. Find the refractive index of the liquid and the diameter of the 5 ^ (th) dark ring.


1
Expert's answer
2021-06-29T10:00:25-0400

"\text {The radius of the k-th bright Newton ring:}"

"r_k=\sqrt{(k-\frac{1}{2})\frac{\lambda R}{n}}(1)"

"\text{measurement results for air:}"

"r_{16}=\frac{1.8}{2}=0.9cm = 0.009m"

"n_E\approx1"

"r_{16}=\sqrt{15.5*\lambda R}"

"\text{measurement results for liquid:}"

"r'_{16}=\frac{1.5}{2}=0.75cm = 0.0075m"

"r'_{16}=\sqrt{15.5*\frac{\lambda R}{n_L}}"


"\frac{r_{16}}{r'_{16}}=\sqrt{n_L}"

"\sqrt{n_L} = \frac{0.009}{0.0075}=1.2"

"n_L=1.44"


"\text{Radius of Newton's k-th dark ring:}"

"r_k=\sqrt{k\frac{\lambda R}{n}}"

"r_5=\sqrt{5*\frac{\lambda R}{n}}"

"\text{From equality (1)}"

"r_{16}=\sqrt{15.5*\frac{\lambda R}{n}}"

"r_5 = r_{16}*\sqrt{\frac{5}{15.5}}=0.005m"

"d_5 = 2*r_5=0.01m"


"\text {Answer: refractive index of the liquid =} 1.44;"

"\text{ the diameter of the 5- th dark ring= }0.01m"